# Enzo Erlich Université Paris Cité, CNRS, IRIF, F-75013, Paris, France and EPITA Research Laboratory (LRE), Paris, France

### v3

@@ 4.2 History-deterministic Reversal-bounded Counter Machines
Let $w=a_{0}\cdots a_{n-1}\in L(\mathcal{A},C)$ with $n>0$, say with accepting run $\rho$. Then, $\rho$ can be decomposed into a sequence $\rho_{0}\cdots\rho_{n-1}$, where each $\rho_{i}$ is an $\varepsilon$-$a_{i}$-$\varepsilon$-path (the exact splits are irrelevant). Consider one of these $\rho_{i}$, say it leads from $q$ to $q^{\prime}$. By construction, $(\mathcal{A}^{\prime},C^{\prime})$ has a $a_{i}$-transition from $q$ to $q^{\prime}$. These transitions form a run $\rho^{\prime}$ of $(\mathcal{A}^{\prime},C^{\prime})$ processing $w$. Note that both runs end in the same state, so $\rho^{\prime}$ ends in an accepting state as well.
@@
Finally, let us remark that HDPA (or equivalently 1-HDRBCM) and deterministic RBCM have incomparable expressiveness. Indeed, the language $E$, which is not accepted by any HDPA (see Theorem 1), can easily be accepted by a deterministic RBCM while the language $N$ (see Example 3) is accepted by an HDPA, but not by any deterministic RBCM [7]. The reason is that these machines are closed under complement, but the complement of $N$ is not accepted by any PA as shown in [7], and therefore also not by any RBCM.
@@ 5 Closure Properties
Fix some ordered alphabet $\Sigma=(a_{0}<a_{1}<\cdots<a_{d-1})$. The Parikh image of a word $w\in\Sigma^{*}$ is the vector $\Phi(w)=(|w|_{a_{0}},|w|_{a_{1}},\ldots,|w|_{a_{d-1}})$ and the Parikh image of a language $L\subseteq\Sigma^{*}$ is $\Phi(L)=\{\Phi(w)\mid w\in L\}$. The commutative closure of $L$ is $\{w\in\Sigma^{*}\mid\Phi(w)\in\Phi(L)\}$.
@@
  • $\Delta=\{((q_{1},q_{2}),(a,(v_{1},v_{2})\cdot\vec{v}_{1}\cdot\vec{v}_{2}),(q_{1}^{\prime},q_{2}^{\prime}))\mid(q_{i},(a,\vec{v}_{i}),q_{i}^{\prime})\in\Delta_{i}\text{ for }i\in\{1,2\}\}$, where $v_{i}$ for $i\in\{1,2\}$ is defined as follows.
    • If $q_{i}$ is the initial state of $\mathcal{A}_{i}$: If $q_{i}^{\prime}\in F_{i}$ then $v_{i}=2$, otherwise $v_{i}=1$.
    • If $q_{i}$ is not the initial state of $\mathcal{A}_{i}$: If $q_{i}\in F_{i}\Leftrightarrow q_{i}^{\prime}\in F_{i}$ then $v_{i}=2$, otherwise $v_{i}=1$.
    Note that this satisfies the following invariant for every nonempty run of $\mathcal{A}$: If $w\in(\Sigma\times D)^{+}$ is the input processed by the run, $\Phi_{e}(w)$ = $(v_{1},v_{2},\ldots)$, and the run ends in state $(q_{1},q_{2})\in Q_{1}\times Q_{2}$, then we have $v_{i}\bmod 2=0$ if and only if $q_{i}\in F_{i}$. Thus, it is the values $v_{i}$ that reflect in the extended Parikh image which of the simulated runs end in an accepting state. Note however, that this only holds for nonempty runs, as we need to initialize the reflection.
@@
$C={}$${}\{n\in\mathbb{N}\mid n\bmod 2=0\}\cdot\mathbb{N}\cdot C_{1}\cdot\mathbb{N}^{d_{2}}\,\cup$
${}\mathbb{N}\cdot\{n\in\mathbb{N}\mid n\bmod 2=0\}\cdot\mathbb{N}^{d_{1}}\cdot C_{2}\cup C_{\varepsilon},$
which is semilinear due to Proposition 1. Then, we have $L(\mathcal{A},C)=L(\mathcal{A}_{1},C_{1})\cup L(\mathcal{A}_{2},C_{2})$ and the following function $r$ is a resolver for $(\mathcal{A},C)$: let $r_{i}(w)=(q_{i},(a,\vec{v}_{i}),q_{i}^{\prime})$ and define $r(w)=((q_{1},q_{2}),(a,(v_{1},v_{2})\cdot\vec{v}_{1}\cdot\vec{v}_{2}),(q_{1}^{\prime},q_{2}^{\prime}))$, where $(v_{1},v_{2})$ is defined as above.
@@
Now, consider closure under inverse homomorphic images. Klaedtke and Rueß have shown that DPA and PA are effectively closed under inverse homomorphic images [29]. We follow their construction to show that HDPA are closed under inverse homomorphic images. In fact, a similar construction is also used to show that regular languages are also closed under inverse homomorphic images [26].
@@
Now, let us consider commutative closure. Cadilhac et al. proved that the commutative closure of any PA (and therefore that of any HDPA) is accepted by some DPA [6], and therefore also by some HDPA.
The negative results follow from a combination of expressiveness results proven in Section 4 and nonexpressiveness results in the literature [6, 7]:
  • Complement: In the first part of the proof of Theorem 1, we show that the language $N$ is accepted by an HDPA, but its complement is known to not be accepted by any PA [7].
  • Concatenation: The language $E$ is the concatenation of the languages $\{a,b\}^{*}$ and $\{a^{n}b^{n}\mid n\in\mathbb{N}\}$, which are both accepted by a DPA, but itself is not accepted by any HDPA (see the proof of Theorem 1).
  • Kleene star: There is a DPA (and therefore also an HDPA) such that the Kleene star of its language is not accepted by any PA [6], and therefore also by no HDPA.
  • Homomorphic image: There is a DPA (and therefore also an HDPA) and a homomorphism such that the homomorphic image of the DPA’s language is not accepted by any PA [6], and therefore also by no HDPA.∎
Table 1 compares the closure properties of HDPA with those of DPA, UCA, and PA. We do not compare to RBCM, as only deterministic ones differ from Parikh automata and results on these are incomplete: However, Ibarra proved closure under union, intersection, and complement [27].
@@
$\cup$ $\cap$ $\overline{\phantom{x}}$cdot $h$ $h^{-1}$ c
@@ 6 Decision Problems
Next, we study various decision problems for history-deterministic PA. First, let us mention that nonemptiness and finiteness are decidable for HDPA, as these problems are decidable for PA [29, 6]. In the following, we consider universality, inclusion, equivalence, regularity, and model checking.
@@
If $w$ has no errors at positions $\{0,1,\ldots,|w|-2\}$, then an induction shows that $(a_{n},c_{0}^{n},c_{1}^{n})$ with
@@
The existence of an error can be captured by a Parikh automaton, leading to the undecidability of the safe word problem for Parikh automata. Let $(\mathcal{A},C)$ be a PA accepting finite words over $\Sigma$. A safe word of $(\mathcal{A},C)$ is an infinite word in $\Sigma^{\omega}$ such that each of its prefixes is in $L(\mathcal{A},C)$.
@@
[…]
@@
Finally, we consider the problems of deciding whether a Parikh automaton is history-deterministic and whether it is equivalent to some HDPA. Both of our proofs follow arguments developed for similar results for history-deterministic pushdown automata [31].
@@
1.) We say that a PA $(\mathcal{A},C)$ is length-complete if for every $n\in\mathbb{N}$, there is some word of length $n$ in $L(\mathcal{A},C)$. Furthermore, let $h\colon\Sigma^{*}\rightarrow\{\#\}^{*}$ be the homomorphism induced by mapping each $a\in\Sigma$ to ##\##. If $(\mathcal{A},C)$ is length-complete, then $\{h(w)\mid w\in L(\mathcal{A},C)\}=\{\#\}^{*}$. Note that the DPA $(\mathcal{A}_{\mathcal{M}},C_{\mathcal{M}})$ we have constructed in the proof of Lemma 3 are length-complete. Thus, we have actually shown that the following problem is undecidable: Given a length-complete DPA $(\mathcal{A},C)$, does it have a safe word?
@@
So, let $(\mathcal{A}^{\prime},C)$ be history-deterministic, i.e., it has a resolver $r\colon\{\#\}^{*}\rightarrow\Delta_{\mathcal{A}^{\prime}}$, where $\Delta_{\mathcal{A}^{\prime}}$ is the set of transitions of $\mathcal{A}^{\prime}$. Note that $r^{*}(\#^{n})$ is a prefix of $r^{*}(\#^{n^{\prime}})$ whenever $n^{\prime}\geqslant n$. Hence, there is a unique limit
@@
Now, as each transition of $\mathcal{A}^{\prime}$ is obtained from a transition of $\mathcal{A}$, there is some infinite word $a_{0}a_{1}a_{2}\cdots\in\Sigma^{\omega}$ such that $(q_{0},(a_{0},\vec{v}_{0}),q_{1})\cdots(q_{n-1},(a_{n-1},\vec{v}_{n-1}),q_{n})$ is an accepting run of $\mathcal{A}$ for every $n\geqslant 1$. Thus, we have $a_{0}\cdots a_{n-1}\in L(\mathcal{A},C)$ for every $n\geqslant 1$. Lastly, we have $\varepsilon\in L(\mathcal{A},C)$, as $\varepsilon\in L(\mathcal{A}^{\prime},C)$. Altogether, $a_{0}a_{1}a_{2}\cdots$ is indeed a safe word for $(\mathcal{A},C)$.
@@
2.) We reduce from the universality problem for PA, which is undecidable [29], via a series of automata transformations.
@@
Finally, let us introduce the model-checking problem (for safety properties): A transition system $\mathcal{T}=(V,v_{I},E,\lambda)$ consists of a finite set $V$ of vertices containing the initial state $v_{I}\in V$, a transition relation $E\subseteq V\times V$, and a labeling function $\lambda\colon V\rightarrow\Sigma$ for some alphabet $\Sigma$. A (finite and initial) path in $\mathcal{T}$ is a sequence $v_{0}v_{1}\cdots v_{n}\in V^{+}$ such that $v_{0}=v_{I}$ and $(v_{i},v_{i+1})\in E$ for all $0\leqslant i<n$. Infinite (initial) paths are defined analogously. The trace of a path $v_{0}v_{1}\cdots v_{n}$ is $\lambda(v_{0})\lambda(v_{1})\cdots\lambda(v_{n})\in\Sigma^{+}$. We denote the set of traces of paths of $\mathcal{T}$ by $\mathsf{tr}(\mathcal{T})$.
Theorem 9.
The model-checking problem for HDPA is decidable.
Let us conclude by mentioning that the dual problem, i.e., given a transition system $\mathcal{T}$ and an HDPA $\mathcal{A}$, does every infinite path of $\mathcal{T}$ have a prefix whose trace is in $L(\mathcal{A})$, is undecidable. This follows from recent results on Parikh automata over infinite words [22], i.e., that model checking for Parikh automata with reachability conditions is undecidable. Such automata are syntactically equal to Parikh automata over finite words and an (infinite) run is accepting if it has a prefix ending in an accepting state whose extended Parikh image is in the semilinear set of the automaton.
@@ 7 Resolvers
So far, a resolver for a HDPA is a function $r\colon\Sigma^{+}\rightarrow\Delta$. This definition does not put any restrictions on the complexity of the function $r$. But how complex is it to resolve the nondeterminism in a HDPA? For history-deterministic finite automata, finite automata with output suffice to resolve their nondeterminism [2] while pushdown automata with output are not sufficient to resolve the nondeterminism in history-deterministic pushdown automata [23, 31].
It is straightforward to show that every HDPA has a positional resolver (i.e., one whose decision is only based on the last state of the run constructed thus far and on the extended Parikh image induced by this run) and that HDPA that have finite-state resolvers (say, implemented by a Mealy machine) can be determinized by taking the product of the HDPA and the Mealy machine. In fact, both proofs are simple adaptions of the corresponding ones for history-deterministic pushdown automata [23, 31]. Hence, we focus on the question whether the nondeterminism in a HDPA can be resolved by functions implemented by DPA.
Fix a PA $(\mathcal{A},C)$ with $\mathcal{A}=(Q,\Sigma\times D,q_{I},\Delta,F)$ throughout this section. We assume w.l.o.g. that $(\mathcal{A},C)$ is complete, i.e., that every letter $a\in\Sigma$ can be processed from every state $q\in Q$. Recall that a function $r\colon\Sigma^{+}\rightarrow\Delta$ is a resolver for $(\mathcal{A},C)$ if, for every $w\in L(\mathcal{A},C)$, we have that $r^{*}(w)$ is an accepting run of $(\mathcal{A},C)$. Here, $r^{*}\colon\Sigma^{*}\rightarrow\Delta^{*}$ is defined as $r^{*}(\varepsilon)=\varepsilon$ and $r^{*}(a_{0}\cdots a_{n})=r^{*}(a_{0}\cdots a_{n-1})\cdot r(a_{0}\cdots a_{n})$. In the following, we call such a resolver a letter-based resolver.
@@
    is also accepted by a $DPA$, as DPA are closed under union. However, by definition of $r$, this is the language of words that do not have a non-Dyck prefix. It is known that this language is not accepted by any PA [7]. Hence, $r$ is not DPA-implementable and, due to uniqueness of $r$, $(\mathcal{A},C)$ does not have DPA-implementable letter-based resolvers.
    Note that $r$ is not even implementable by PA (with the obvious definition), as the language of words without non-Dyck prefixes is not even accepted by (possibly nondeterministic) PA.
Note that a transition-based resolver operates on the run constructed thus far, which contains information on whether there was a non-Dyck prefix in the word processed thus far: If there is one, then the resolver continued the run on that prefix by moving to state $q_{n}$. Thus, any run consistent with the resolver ends in state $q_{n}$ if and only if the word processed thus far has a non-Dyck prefix. This saves a transition-based resolver from checking every input for non-Dyckness, a property that cannot be checked by an (even nondeterministic) PA. A letter-based resolver, on the other hand, does not get the information in which state the run constructed thus far ends, and a DPA-implementable one can also not compute that information. Thus, DPA-implementable letter-based resolvers are strictly weaker than DPA-implementable transition-based resolvers.
@@
Fix a PA $(\mathcal{A},C)$ with $\mathcal{A}=(Q,\Sigma\times D,q_{I},\Delta,F)$. We assume (w.l.o.g.) that $(\mathcal{A},C)$ is complete in the following sense: for every $q\in Q$ and every $a\in\Sigma$, there is a $\vec{v}\in D$ and a $q^{\prime}\in Q$ such that $(q,(a,\vec{v}),q^{\prime})\in\Delta$, i.e., from every state, every letter can be processed. As usual, an incomplete automaton can be completed by adding a fresh non-accepting sink state.
We define the one-token game on $(\mathcal{A},C)$, a perfect-information zero-sum game played between two players called Challenger and Resolver. The positions of the game consist of pairs $(q,q^{\prime})\in Q\times Q$. The game is played in rounds $i=0,1,2,\ldots$ starting from the initial position $(q_{0},q_{0}^{\prime})=(q_{I},q_{I})$. In round $i$ at position $(q_{i},q_{i}^{\prime})$:
@@
Resolver wins the one-token game on a PA $(\mathcal{A},C)$ if and only if $(\mathcal{A},C)$ is a HDPA.
@@
The proof here is analogous to the one for finite automata [33], but we need to repeat it here as our further results depend on properties of the one-token game that are shown in this proof.
@@
Now, we define the function $t\colon P(\mathcal{A},C)\times\Sigma\rightarrow\Delta$. Let $\rho\in P(\mathcal{A},C)$ be a run prefix of $(\mathcal{A},C)$ processing the word $w$, let $a\in\Sigma$, and let $\delta$ be the transition picked by $\sigma$ when Challenger has picked $w$ thus far, both players have picked $\rho$, and Challenger has picked $a$ to start the next round. Then, we define $t(\rho,a)=\delta$. Note that $\delta$ must be residual w.r.t. $\rho$ and $a$.
@@
Let $W$ be the winning region of Resolver in the safety game and consider a vertex of the form $(q,\vec{v},q,\vec{v},a)\in W$, i.e., it is Resolver’s turn. Then, there is at least one successor in the winning region $W$, which must be of the form $(q,\vec{v},q,\vec{v},a,\delta)$ for a residual transition $\delta$, as argued above. As argued above, Resolver can pick any residual transition to win the one-token game. Hence, we will construct a DPA-implementable resolver that will always pick a residual transition.
However, in general there might be several residual transitions available. So, to make define the resolver formally, we need to select one such transition. To this end, let $\delta_{0}<\delta_{1}<\cdots\delta_{n}$ be a fixed ordering of the set $\Delta$ of transitions of $(\mathcal{A},C)$. We fix a letter $a$ and a transition $\delta_{i}$ and construct a DPA that does the following, when given a finite run $\rho$:
@@
Note that if the inherent structure of the safety games we consider here suffices to show that the winning region is indeed semilinear, then the DPA implement a resolver cannot be effectively constructed, as it is undecidable whether a PA is history-deterministic (Theorem 8).

### v4

@@ 4.2 History-deterministic Reversal-bounded Counter Machines
Let $w=a_{0}\cdots a_{n-1}\in L(\mathcal{A},C)$ with $n>0$, say with accepting run $\rho$. Then, $\rho$ can be decomposed into a sequence $\rho_{0}\cdots\rho_{n-1}$, where each $\rho_{i}$ is an $\varepsilon$-$a_{i}$-$\varepsilon$-path (the exact splits are irrelevant). Consider one of these $\rho_{i}$, say it leads from $q$ to $q^{\prime}$. By construction, $(\mathcal{A}^{\prime},C^{\prime})$ has an $a_{i}$-transition from $q$ to $q^{\prime}$. These transitions form a run $\rho^{\prime}$ of $(\mathcal{A}^{\prime},C^{\prime})$ processing $w$. Note that both runs end in the same state, so $\rho^{\prime}$ ends in an accepting state as well.
@@
Finally, let us remark that HDPA (or equivalently 1-HDRBCM) and deterministic RBCM have incomparable expressiveness. Indeed, the language $E$, which is not accepted by any HDPA (see Theorem 1), can easily be accepted by a deterministic RBCM while the language $N$ (see Example 3) is accepted by an HDPA, but not by any deterministic RBCM. The reason is that these machines are closed under complement, but the complement of $N$ is not accepted by any PA [9, Proposition 11], and therefore also not by any RBCM.
@@ 5 Closure Properties
Fix some alphabet $\Sigma=\{a_{0},a_{1},\cdots,a_{d-1}\}$ with fixed order $a_{0}<a_{1}<\cdots<a_{d-1}$. The Parikh image of a word $w\in\Sigma^{*}$ is the vector $\Phi(w)=(|w|_{a_{0}},|w|_{a_{1}},\ldots,|w|_{a_{d-1}})$ and the Parikh image of a language $L\subseteq\Sigma^{*}$ is $\Phi(L)=\{\Phi(w)\mid w\in L\}$. The commutative closure of $L$ is $\{w\in\Sigma^{*}\mid\Phi(w)\in\Phi(L)\}$.
@@
  • $\Delta=\{((q_{1},q_{2}),(a,(f_{1},f_{2})\cdot\vec{v}_{1}\cdot\vec{v}_{2}),(q_{1}^{\prime},q_{2}^{\prime}))\mid(q_{i},(a,\vec{v}_{i}),q_{i}^{\prime})\in\Delta_{i}\text{ for }i\in\{1,2\}\}$, where $f_{i}$ for $i\in\{1,2\}$ is defined as follows.
    • If $q_{i}$ is the initial state of $\mathcal{A}_{i}$: If $q_{i}^{\prime}\in F_{i}$ then $f_{i}=2$, otherwise $f_{i}=1$.
    • If $q_{i}$ is not the initial state of $\mathcal{A}_{i}$: If $q_{i}\in F_{i}\Leftrightarrow q_{i}^{\prime}\in F_{i}$ then $f_{i}=2$, otherwise $f_{i}=1$.
    Note that this satisfies the following invariant for every nonempty run of $\mathcal{A}$: If $w\in(\Sigma\times D)^{+}$ is the input processed by the run, $\Phi_{e}(w)$ = $(f_{1},f_{2},\ldots)$, and the run ends in state $(q_{1},q_{2})\in Q_{1}\times Q_{2}$, then we have $f_{i}\bmod 2=0$ if and only if $q_{i}\in F_{i}$. Thus, it is the values $f_{i}$ that reflect in the extended Parikh image which of the simulated runs end in an accepting state. Note however, that this only holds for nonempty runs, as we need to initialize the reflection.
@@
$C=\{n\in\mathbb{N}\mid n\bmod 2=0\}\cdot\mathbb{N}\cdot C_{1}\cdot\mathbb{N}^{d_{2}}\,\cup\mathbb{N}\cdot\{n\in\mathbb{N}\mid n\bmod 2=0\}\cdot\mathbb{N}^{d_{1}}\cdot C_{2}\cup C_{\varepsilon},$
which is semilinear due to Proposition 1. Then, we have $L(\mathcal{A},C)=L(\mathcal{A}_{1},C_{1})\cup L(\mathcal{A}_{2},C_{2})$ and the following function $r$ is a resolver for $(\mathcal{A},C)$: let $r_{i}(w)=(q_{i},(a,\vec{v}_{i}),q_{i}^{\prime})$ and define $r(w)=((q_{1},q_{2}),(a,(f_{1},f_{2})\cdot\vec{v}_{1}\cdot\vec{v}_{2}),(q_{1}^{\prime},q_{2}^{\prime}))$, where $(f_{1},f_{2})$ is defined as above.
@@
Now, consider closure under inverse homomorphic images. Klaedtke and Rueß have shown that DPA and PA are effectively closed under inverse homomorphic images [32, Property 4]. We follow their construction to show that HDPA are closed under inverse homomorphic images. In fact, a similar construction is also used to show that regular languages are also closed under inverse homomorphic images [29].
@@
Now, let us consider commutative closure. Cadilhac et al. proved that the commutative closure of any PA (and therefore that of any HDPA) is accepted by some DPA [8, Proposition 3.17], and therefore also by some HDPA.
The negative results follow from a combination of expressiveness results proven in Section 4 and nonexpressiveness results in the literature:
  • Complement: In the first part of the proof of Theorem 1, we show that the language $N$ is accepted by an HDPA, but its complement is known to not be accepted by any PA [9, Proposition 11].
  • Concatenation: The language $E$ is the concatenation of the languages $\{a,b\}^{*}$ and $\{a^{n}b^{n}\mid n\in\mathbb{N}\}$, which are both accepted by a DPA (and therefore also by HDPA), but itself is not accepted by any HDPA (see the proof of Theorem 1).
  • Kleene star: There is a DPA (and therefore also an HDPA) such that the Kleene star of its language is not accepted by any PA [8, Proposition 3.17], and therefore also by no HDPA.
  • Homomorphic image: The language
    $\{a,b\}^{*}\cdot\{ca^{n-1}b^{n}\mid n>0\}$
    is accepted by the DPA $(\mathcal{A},C)$ with $\mathcal{A}$ as in Figure 7 and $C=\{(n,n)\mid n>0\}$, and thus also by an HDPA.
    $a,(0,0)$$b,(0,0)$$c,(1,0)$$a,(1,0)$$b,(0,1)$$b,(0,1)$
    Figure 7: The automaton for the language $\{a,b\}^{*}\cdot\{ca^{n-1}b^{n}\mid n>0\}$.
    However, its homomorphic image under the morphism $h$ uniquely identified by $h(a)=h(c)=a$ and $h(b)=b$ is $E$, which is not accepted by any HDPA, as shown in Theorem 1.∎
Table 1 compares the closure properties of HDPA with those of DPA, UCA, and PA. We do not compare to RBCM, as only deterministic ones differ from Parikh automata and results on these are incomplete: However, Ibarra proved closure under union, intersection, and complement [30, Lemma 3.1 and Lemma 3.2].
@@
union intersection complement concate- Kleene homo- inverse commutative
nation star morphism homomorphism closure
@@ 6 Decision Problems
Next, we study various decision problems for history-deterministic PA. First, let us mention that nonemptiness and finiteness are decidable for HDPA, as these problems are decidable for PA [32, Property 6][8, Proposition 3.16]. In the following, we consider universality, inclusion, equivalence, regularity, and model checking.
@@
If $w$ has no errors at positions $\{0,1,\ldots,|w|-2\}$, then an induction over $n\in\mathbb{N}$ shows that $(a_{n},c_{0}^{n},c_{1}^{n})$ with
@@
The existence of an error can be captured by a Parikh automaton, leading to the undecidability of the safe word problem for Parikh automata. Let $(\mathcal{A},C)$ be a PA accepting finite words over $\Sigma$. A safe word of $(\mathcal{A},C)$ is an infinite word over $\Sigma$ such that each of its prefixes is in $L(\mathcal{A},C)$.
@@
Next, let us introduce the model-checking problem (for safety properties): A transition system $\mathcal{T}=(V,v_{I},E,\lambda)$ consists of a finite set $V$ of vertices containing the initial state $v_{I}\in V$, a transition relation $E\subseteq V\times V$, and a labeling function $\lambda\colon V\rightarrow\Sigma$ for some alphabet $\Sigma$. A (finite and initial) path in $\mathcal{T}$ is a sequence $v_{0}v_{1}\cdots v_{n}\in V^{+}$ such that $v_{0}=v_{I}$ and $(v_{i},v_{i+1})\in E$ for all $0\leqslant i<n$. Infinite (initial) paths are defined analogously. The trace of a path $v_{0}v_{1}\cdots v_{n}$ is $\lambda(v_{0})\lambda(v_{1})\cdots\lambda(v_{n})\in\Sigma^{+}$. We denote the set of traces of paths of $\mathcal{T}$ by $\mathsf{tr}(\mathcal{T})$.
Theorem 8.
The model-checking problem for HDPA is decidable.
Let us conclude by mentioning that the dual problem, i.e., given a transition system $\mathcal{T}$ and an HDPA $\mathcal{A}$, does every infinite path of $\mathcal{T}$ have a prefix whose trace is in $L(\mathcal{A})$, is undecidable. This follows from recent results on Parikh automata over infinite words [25, Theorem 17], i.e., that model checking for Parikh automata with reachability conditions is undecidable. Such automata are syntactically equal to Parikh automata over finite words and an (infinite) run is accepting if it has a prefix ending in an accepting state whose extended Parikh image is in the semilinear set of the automaton.
@@
Nonemptiness Finiteness Universality Inclusion Equivalence Regularity Model Checking
@@
Finally, we consider the problems of deciding whether a Parikh automaton is history-deterministic and whether it is equivalent to some HDPA. Both of our proofs follow arguments developed for similar results for history-deterministic pushdown automata [34, Theorem 6.1].
@@
1.) We say that a PA $(\mathcal{A},C)$ is length-complete if for every $n\in\mathbb{N}$, there is some word of length $n$ in $L(\mathcal{A},C)$. Furthermore, let $h\colon\Sigma^{*}\rightarrow\{\#\}^{*}$ be the homomorphism induced by mapping each $a\in\Sigma$ to ##\##. If $(\mathcal{A},C)$ is length-complete, then $\{h(w)\mid w\in L(\mathcal{A},C)\}=\{\#\}^{*}$. Note that the DPA’s $(\mathcal{A}_{\mathcal{M}},C_{\mathcal{M}})$ we have constructed in the proof of Lemma 3 are length-complete. Thus, we have actually shown that the following problem is undecidable: Given a length-complete DPA $(\mathcal{A},C)$, does it have a safe word?
@@
So, let $(\mathcal{A}^{\prime},C)$ be history-deterministic, i.e., it has a resolver $r\colon\{\#\}^{*}\rightarrow\Delta_{\mathcal{A}^{\prime}}$, where $\Delta_{\mathcal{A}^{\prime}}$ is the set of transitions of $\mathcal{A}^{\prime}$. Note that $r^{*}(\#^{n})$ is a prefix of $r^{*}(\#^{n^{\prime}})$ whenever $n^{\prime}\geqslant n$. Hence, there is a unique infinite sequence
@@
Now, as each transition of $\mathcal{A}^{\prime}$ is obtained from a transition of $\mathcal{A}$, there is some infinite word $a_{0}a_{1}a_{2}\cdots$ over $\Sigma$ such that $(q_{0},(a_{0},\vec{v}_{0}),q_{1})\cdots(q_{n-1},(a_{n-1},\vec{v}_{n-1}),q_{n})$ is an accepting run of $\mathcal{A}$ for every $n\geqslant 1$. Thus, we have $a_{0}\cdots a_{n-1}\in L(\mathcal{A},C)$ for every $n\geqslant 1$. Lastly, we have $\varepsilon\in L(\mathcal{A},C)$, as $\varepsilon\in L(\mathcal{A}^{\prime},C)$. Altogether, $a_{0}a_{1}a_{2}\cdots$ is indeed a safe word for $(\mathcal{A},C)$.
@@
2.) We reduce from the universality problem for PA, which is undecidable [32, Property 7], via a series of automata transformations.
@@ 7 Resolvers
So far, a resolver for an HDPA is a function $r\colon\Sigma^{+}\rightarrow\Delta$. This definition does not put any restrictions on the complexity of the function $r$. But how complex is it to resolve the nondeterminism in an HDPA? For history-deterministic finite automata, finite automata with output suffice to resolve their nondeterminism [3, Lemma 1] while pushdown automata with output are not sufficient to resolve the nondeterminism in history-deterministic pushdown automata [34, Theorem 7.3][26, Lemma 10.2].
It is straightforward to show that every HDPA has a positional resolver (i.e., one whose decision is only based on the last state of the run constructed thus far and on the extended Parikh image induced by this run) and that HDPA that have finite-state resolvers (say, implemented by a Mealy machine) can be determinized by taking the product of the HDPA and the Mealy machine. In fact, both proofs are simple adaptions of the corresponding ones for history-deterministic pushdown automata [26, Lemma 10.1 and Remark 10.3]. Hence, we focus on the question whether the nondeterminism in an HDPA can be resolved by functions implemented by more expressive formalisms, i.e., we are looking for deterministic machines that implement resolvers for HDPA. A natural question is then whether DPA are sufficient for this task, i.e., the question whether the nondeterminism in an HDPA can be resolved by DPA.
Let us formalize this question. Fix a PA $(\mathcal{A},C)$ with $\mathcal{A}=(Q,\Sigma\times D,q_{I},\Delta,F)$ throughout this section. We assume (without loss of generality) that $(\mathcal{A},C)$ is complete. Recall that a function $r\colon\Sigma^{+}\rightarrow\Delta$ is a resolver for $(\mathcal{A},C)$ if, for every $w\in L(\mathcal{A},C)$, we have that $r^{*}(w)$ is an accepting run of $(\mathcal{A},C)$. Here, $r^{*}\colon\Sigma^{*}\rightarrow\Delta^{*}$ is defined as $r^{*}(\varepsilon)=\varepsilon$ and $r^{*}(a_{0}\cdots a_{n})=r^{*}(a_{0}\cdots a_{n-1})\cdot r(a_{0}\cdots a_{n})$. In the following, we call such a resolver a letter-based resolver.
@@
    is also accepted by a $DPA$, as DPA are closed under union. However, by definition of $r$, this is the language of words that do not have a non-Dyck prefix. It is known that this language is not accepted by any PA [9, Proposition 11]. Hence, $r$ is not DPA-implementable and, due to uniqueness of $r$, $(\mathcal{A},C)$ does not have DPA-implementable letter-based resolvers.
    Note that $r$ is not even implementable by PA (with the obvious definition), as the language of words without non-Dyck prefixes is not even accepted by (possibly nondeterministic) PA [9, Proposition 11].
Note that a transition-based resolver operates on the run constructed thus far, which contains information on whether there was a non-Dyck prefix in the word processed thus far: If there is one, then the resolver continued the run on that prefix by moving to state $q_{n}$. Thus, any run consistent with the resolver ends in state $q_{n}$ if and only if the word processed thus far has a non-Dyck prefix. This saves a transition-based resolver from checking every input for non-Dyckness, a property that cannot be checked by an (even nondeterministic) PA. A letter-based resolver, on the other hand, does not get the information in which state the run constructed thus far ends, and a DPA-implementable one also cannot compute that information. Thus, DPA-implementable letter-based resolvers are strictly weaker than DPA-implementable transition-based resolvers.
@@
Recall that we fixed a complete PA $(\mathcal{A},C)$ with $\mathcal{A}=(Q,\Sigma\times D,q_{I},\Delta,F)$. We define the one-token game on $(\mathcal{A},C)$, a perfect-information zero-sum game played between two players called Challenger and Resolver. The positions of the game consist of pairs $(q,q^{\prime})\in Q\times Q$. The game is played in rounds $i=0,1,2,\ldots$ starting from the initial position $(q_{0},q_{0}^{\prime})=(q_{I},q_{I})$. In round $i$ at position $(q_{i},q_{i}^{\prime})$:
@@
Resolver wins the one-token game on a PA $(\mathcal{A},C)$ if and only if $(\mathcal{A},C)$ is an HDPA.
@@
The proof here is analogous to the one for finite automata [36, Lemma 7 (case of $k=1$)], but we need to repeat it here as our further results depend on properties of the one-token game that are shown in this proof.
@@
We define the function $t\colon P(\mathcal{A},C)\times\Sigma\rightarrow\Delta$. Let $\rho\in P(\mathcal{A},C)$ be a run prefix of $(\mathcal{A},C)$ processing the word $w$, let $a\in\Sigma$, and let $\delta$ be the transition picked by $\sigma$ when Challenger has picked $w$ thus far, both players have picked $\rho$, and Challenger has picked $a$ to start the next round. Then, we define $t(\rho,a)=\delta$. Note that $\delta$ must be residual w.r.t. $\rho$ and $a$, as argued above.
@@
Let $W$ be the winning region of Resolver in the safety game and consider a vertex of the form $(q,\vec{v},q,\vec{v},a)\in W$, i.e., it is Resolver’s turn. Then, there is at least one successor in the winning region $W$, which must be of the form $(q,\vec{v},q,\vec{v},a,\delta)$ for a residual transition $\delta$, as argued above. Thus, Resolver can pick any residual transition to win the one-token game, again as argued above. Hence, we will construct a DPA-implementable resolver that will always pick a residual transition.
However, in general there might be several residual transitions available. So, to define the resolver formally, we need to select one such transition. To this end, let $\delta_{0}<\delta_{1}<\cdots\delta_{n}$ be a fixed ordering of the set $\Delta$ of transitions of $(\mathcal{A},C)$. We fix a letter $a$ and a transition $\delta_{i}$ and construct a DPA that does the following, when given a finite run $\rho$:
@@
Note that if the inherent structure of the safety games we consider here suffices to show that the winning region is indeed semilinear, then the DPA’s implementing a resolver cannot be effectively constructed, as it is undecidable whether a PA is history-deterministic (Theorem 9).